Show that on the curve, y= 3x^2-7x+6, the chord joining the points whose abscissa are x=1 and x=2, parallel to the tangent at the point whose abscissa is x=3/2


1
Expert's answer
2021-05-14T05:45:56-0400

Solution

Ordinates for two points of chord abscissa x1=1 and x2=2 are y1=y(x1)=2 and y2= y(x2)=4.

The slope of the chord is k=( y2-y1)/( x2-x1)=2.

The tangent at the any point of curve y(x) is the line with the slope equal to its derivative y’(x)=6x-7. At the abscissa x0=3/2 the tangent is y’(3/2)=2.

So slopes of the chord and the tangent are equal. Therefore this lines are parallel.